OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y=sin(sinx) (π,0)

zepdrix (zepdrix):

\[\Large\rm y=\sin(\sin x),\qquad (\pi,0)\]

zepdrix (zepdrix):

Were you able to find a derivative of your function y?

OpenStudy (anonymous):

would it be cos^2x(sinx)?

zepdrix (zepdrix):

No.\[\Large\rm y'=\cos(\sin x)\cdot \color{royalblue}{\left[\sin x\right]'}\]\[\Large\rm y'=\cos x~ \cos(\sin x)\]Bit of a tricky one to differentiate :) Understand how I chain ruled there?

OpenStudy (anonymous):

yeah i understand how now